In triangle ABC if medianAD makes an angle they with BC and AB=2AD then sin theta

Grade:9

1 Answers

Vikas TU
4 years ago
Dear student 
Let median be x = AD 
Then AB = 2X 
DC / SinQ = AD / SinC 
a/2 /SinQ = x/ Sin c
In triangle ABC 
AB /SinC = BC /SinA 
2x/SinC =  a/SinA 
x/SinC = a/2SinB = a/2 / SinQ  
a/ 2SinA = a/2sinQ 
Hence SinQ = SinA